Vandalay Industries is considering the purchase of a new machine for the production of latex. Machine A costs $2,060,000 and will last for 6 years. Variable costs are 37 percent of sales, and fixed costs are $166,000 per year. Machine B costs $4,220,000 and will last for 9 years. Variable costs for this machine are 32 percent of sales and fixed costs are $84,000 per year. The sales for each machine will be $8.44 million per year. The required return is 10 percent and the tax rate is 21 percent. Both machines will be depreciated on a straight-line basis. If the company plans to replace the machine when it wears out on a perpetual basis, what is the EAC for machine A? If the company plans to replace the machine when it wears out on a perpetual basis, what is the EAC for machine B?
Answer #1

NPV for Machine A:

  • Fixed Investment = $2,060,000 for 6 years. Using straight line method, Depreciation = $2,060,000 / 6 = $343,333.33
  • Sales = $8,440,000
  • Cost = Variable cost + Fixed cost = 37% * Sales + Fixed cost = 0.37 * $8,440,000 + $166,000 = $3,288,800

Year-wise Incremental Cash Flow, ICF = Operating cash flow + Depreciation

  • Operating cash flow, OCF = (Sales - cost - depreciation) * (1-tax rate)
  • Depreciation
  • ICF = OCF + Depreciation
    • For the first year, OCF = ($8,440,000 - $3,288,800 - $343,333.33 ) * (1-0.21) = $3,798,214.67
    • Depreciation = $343,333.33
    • ICF = $3,798,214.67 +  $343,333.33 = $4,141,548

Year-wise ICF:

Year                              1                           2                           3                           4                           5                           6
ICF       4,141,548.00    4,141,548.00    4,141,548.00    4,141,548.00    4,141,548.00    4,141,548.00

PV(ICF) = $4,141,548/(1+0.10) + $4,141,548/(1+0.10)^2 + ..... + $4,141,548/(1+0.10)^6 = $18,037,521.24

EAC = NPV/Annuity factor

  • NPV = Initial outlay + PV(ICF) = -$2,060,000 + $18,037,521.24 = $15,977.521.24
  • Annuity factor = (1+0.10)^6 = 1.771561
  • EAC = $15,977.521.24 / 1.771561
  • EAC = $9,018,894.21

Machine B:

  • Fixed Investment = $4,220,000 for 9 years. Depreciation = $4,220,000 / 9 = $468,888.89
  • Sales = $8,440,000
  • Cost = Variable cost + Fixed cost = 32% * Sales + Fixed cost = 0.32 * $8,440,000 + $84,000 = $2,784,800

ICF = OCF + Depreciation

  • For first year, OCF = ($8,440,000 - $2,784,800 - $468,888.89 ) * (1-0.21) = $4,097,185.78
  • Depreciation = $468,888.89
  • ICF = $4,097,185.78 +  $468,888.89 = $4,566,074.67

Year-wise ICF:

Year 1 2 3 4 5 6 7 8 9
ICF       4,566,074.67    4,566,074.67    4,566,074.67    4,566,074.67    4,566,074.67    4,566,074.67    4,566,074.67    4,566,074.67    4,566,074.67

PV(CF) = $4,566,074.67/(1+0.10) + $4,566,074.67/(1+0.10)^2 + ..... + $4,566,074.67/(1+0.10)^9 = $26,296,132.77

EAC = NPV/Annuity factor

  • NPV = - $4,220,000 + $26,296,132.77= $22,076,132.77
  • Annuity factor = (1+0.10)^9 = 2.357948
  • EAC = $22,076,132.77/ 2.357948
  • EAC = $9,362,435.33